Find the angles which is equal to its supplement
 Find the angles which is equal to its supplement

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the measure of the required angle be x^o .
We know that, sum of measures of supplementary angle pair is 180 .
Then,
= x + x = 180^o
= 2x = 180^o
= x = 180/2
= x = 90^o
Hence, the required angle measures is 90^o

Thanks